Exercise 6-12 Variable Costing Income Statement; Reconciliation [LO6-2, LO6-3]

Whitman Company has just completed its first year of operations. The company’s absorption costing income statement for the year follows:

Whitman Company
Income Statement
Sales (38,000 units × $43.60 per unit) $ 1,656,800
Cost of goods sold (38,000 units × $21 per unit) 798,000
Gross margin 858,800
Selling and administrative expenses 437,000
Net operating income $ 421,800

The company’s selling and administrative expenses consist of $285,000 per year in fixed expenses and $4 per unit sold in variable expenses. The $21 unit product cost given above is computed as follows:

Direct materials $ 10
Direct labor 3
Variable manufacturing overhead 3
Fixed manufacturing overhead ($270,000 ÷ 54,000 units) 5
Absorption costing unit product cost $ 21

Required:

1. Redo the company’s income statement in the contribution format using variable costing.

2. Reconcile any difference between the net operating income on your variable costing income statement and the net operating income on the absorption costing income statement above.
Answer #1
1
Variable Costing Income Statement
Sales 1656800
Variable expenses:
Variable cost of goods sold 608000 =38000*(10+3+3)
Variable selling and administrative expenses 152000 =38000*4
760000
Contribution margin 896800
Fixed expenses:
Fixed manufacturing overhead 270000
Fixed selling and administrative expenses 285000
555000
Net operating income 341800
2
Variable costing net operating income (loss) 341800
Add: Fixed manufacturing overhead cost
deferred in inventory under absorption costing
80000 =(54000-38000)*5
Absorption costing net operating income (loss) 421800